A 0.371-M aqueous solution of C5H5N (pyridine) has a pH of 9.37. Calculate the pH of a buffer solution that is 0.371 M in C5H5N and 0.159 M in C5H5NH+.

pH =
Answer #1

First, identify the pKb from the intitial concentration:

B+ H"O <-< HB+ + OH-

Kb = [BH+][OH-]/[B]

pOH = 14-pH = 14-9.37 = 4.63

[OH-] = 10^-pOH = 10^-4.63 =2.34410^-5

[BH+] = [OH-] =2.34410^-5

[B] = 0.371-2.34410^-5 = 0.3568707 M

Kb = (2.34410^-5)(2.34410^-5) /(0.3568707) = 0.00055940682 = 5.59*10^-4

now

pKb = -log(Kb) = -log(5.59*10^-4) = 3.2525

p

for the buffer

pOH = pKB + log(C5H5NH+/C5H5N)

pOH = 3.2525+ log(0.159/0.371) = 2.8845

pH = 14-2.8845 = 11.1155